write the equation of the line that passes through the two points.Use integer coefficients (9,-14)and (3,-10). Please also graph the line.


Answer by mananth(16946)

x1 y1 x2 y2
9 -14 3 -10

slope m = (y2-y1)/(x2-x1)
( -10 - -14 )/( 3 - 9 )
( 4 / -6 )
m= - 2/ 3
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( 8 , 1 ) in
Y = m x + b
1.00 = - 2/3 * 8 + b
1.00 = -16/3 + b
b= 1 + 16/3
b= 19/ 3
So the equation will be
Y = - 2/3 x + 19/3
Generate values of y from x
x=2, y = 5 (2,5)
x=5,y=3 (5,3)
x=-4,9 (-4,9)